OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.B(6, 2), m = @amistre64

OpenStudy (anonymous):

M= -.05 @Catlover5925

OpenStudy (catlover5925):

@sammixboo plz help i am out of it right now

OpenStudy (amistre64):

we need to recall what standard form of a line looks like. then we can approach a solution

OpenStudy (anonymous):

Y=mx+b

OpenStudy (amistre64):

y = mx + b is called slope intercept form, not standard form standard from looks like: Ax + By = C, and A is positive.

OpenStudy (amistre64):

now, what we are given is a point, and a slope; so we can build a point slope form and then math it into a standard form

OpenStudy (anonymous):

how do you get point slope form

OpenStudy (amistre64):

given a point (xo,yo) and a slope, a/b, the point slope format is: y-yo = a/b (x-xo) expanding it out and getting the parts in the right places is all thats left y - yo = a/b x - a/b xo -a/b x + y = yo -a/b xo -ax + by = b yo -a xo ax - by = a xo - b yo

OpenStudy (amistre64):

-.05 = -5/100 = -1/20 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then a=-1, and b=20, xo=6, yo=2

OpenStudy (amistre64):

since we have a -a, we might want to just use the next to last setup -ax + by = b yo -a xo

OpenStudy (amistre64):

or let b=-20 either way

OpenStudy (amistre64):

ax - by = a xo - b yo 1x - -20y = 1(6) - -20(2) x +20y = 46

OpenStudy (anonymous):

Thanks so much

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

@amistre64 The answer has _+_____ as its slots.

OpenStudy (amistre64):

not sure what you mean by that .... a picture may be more understandbale

OpenStudy (anonymous):



OpenStudy (anonymous):

The underlined spaces are the ones that need to be filled except for the one underneath the + sign. @amistre64

OpenStudy (amistre64):

then im sure this will fit x + 20y = 46

OpenStudy (anonymous):

It will not let me put the six on the end

OpenStudy (catlover5925):